# Functional-level Uncertainty Quantification for Calibrated Fine-tuning on LLMs

### v1

@@ Abstract
  1. From common-sense reasoning to domain-specific tasks, parameter-efficient fine tuning (PEFT) methods for large language models (LLMs) have showcased significant performance improvements on downstream tasks. However, fine-tuned LLMs often struggle with overconfidence in uncertain predictions, particularly due to sparse training data. This overconfidence reflects poor epistemic uncertainty calibration, which arises from limitations in the model’s ability to generalize with limited data. Existing PEFT uncertainty quantification methods for LLMs focus on the post fine-tuning stage and thus have limited capability in calibrating epistemic uncertainty. To address these limitations, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which captures and calibrates functional-level epistemic uncertainty during the fine-tuning stage via a mixture-of-expert framework. We show that UQ4CT reduces Expected Calibration Error (ECE) by more than $25\%$ while maintaining high accuracy across $5$ benchmarks. Furthermore, UQ4CT maintains superior ECE performance with high accuracy under distribution shift, showcasing improved generalizability.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized various domains as general task solvers (Chang et al., 2024). To adapt LLMs for specific downstream tasks or create instruction-following models, fine-tuning have become increasingly important (Houlsby et al., 2019; Hu et al., 2021a; Liu et al., 2022; Ding et al., 2022, 2023). This involves additional training on pre-trained LLMs using a smaller dataset (Zhong et al., 2021; Ren et al., 2022). Through fine-tuning, the model parameters are updated to better adapt to the domain-specific knowledge (Peng et al., 2023). To reduce the computational cost for fine-tuning, Hu et al. (2021a) proposed Low-Rank Adaptation (LoRA), which effectively reduces the parameters required for fine-tuning by introducing low-rank trainable matrices at each layer of the transformer architecture instead of fine-tuning the full model parameters. Li et al. (2024); Wu et al. (2024a) proposed LoRA Mixture-of-Experts (MoE) models which grants better performance while maintaining parameter efficiency.
However, previous studies have shown that fine-tuned LLMs are often overconfident with their predictions (Xiao et al., 2022a; He et al., 2023; Tian et al., 2023; OpenAI, 2023). This resembles poorly calibrated uncertainty (Zhou et al., 2022) due to the sparsity of fine-tuning data. Overconfidence is a crucial problem in safety-related decision making or in fields where data is very limited, such as experimental design, climate science and medical diagnosis (Singhal et al., 2022; Wu et al., 2023a; Lampinen et al., 2023; Li et al., 2022). Thus, methods that enhance uncertainty quantification of fine-tuned LLMs is urgently needed to assure trustworthy predictions for better application.
Established uncertainty quantification methods have been studied in conjunction with the LoRA structure. Monte-Carlo dropout (Gal and Ghahramani, 2016a) interprets dropout in neural networks as approximate Bayesian inference in deep Gaussian processes, allowing uncertainty estimates to be obtained from existing LoRA adapters without modifying them. Checkpoint ensemble (Chen et al., 2017) utilizes predictions from multiple LoRA checkpoints saved during a single fine-tuning process to calibrate uncertainty. Deep ensemble (Lakshminarayanan et al., 2017; Wang et al., 2023; Zhai et al., 2023a) combines the predictions from multiple LoRA adapters for better uncertainty calibration. Laplace-LoRA (Yang et al., 2024a) applies Bayesian inference via Laplace approximation to the LoRA parameters after fine-tuning, resulting in improved calibration and uncertainty estimates. Although these methods have demonstrated improved uncertainty estimations, they all utilize a fixed set of LoRA parameters fine-tuned over the entire downstream task dataset. The point estimates of parameters have very limited capabilities in capturing epistemic uncertainty, while direct calibration of epistemic uncertainty over the entire LoRA parameter space is an ideal but not practical approach.
@@
Therefore, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT) to calibrate the functional-level epistemic uncertainty via the LoRA MoE architecture. We propose a functional perspective on LoRA parameters where we treat the LoRA experts as basis functions and consider the more complex, prompt dependent functions as mixtures of those basis functions. On top of learning the parameters in the LoRA experts, UQ4CT also trains a prompt-dependent LoRA mixture to form a calibrated distribution over the functional space. The LoRA experts capture different functional relationships in the fine-tuning data throughout training, and the MoE routers dynamically select these functional bases conditioned on the input. The selection process models the functional level epistemic uncertainty, and consequently captures the uncertainty in the output space. We calibrate functional level epistemic uncertainty to align with predictive correctness during training time. This significantly improves uncertainty estimations of the model on its predictions without compromising the accuracy. To summarize, our contributions include:
@@
  • A novel epistemic uncertainty quantification approach with Mixture-of-Experts architecture during fine-tuning stage to model functional level epistemic uncertainty and align with predictive correctness, which mitigates overconfidence issue and improves generalizability.
@@
  • A novel training calibration loss function incorporating predictive correctness to calibrate the prompt-dependent LoRA mixture for better uncertainty estimation.
@@
  • More than $25\%$ Expected Calibration Error reduction on $4$ common-sense reasoning tasks and $1$ domain-specific question answering task, superior ECE performance under distribution shift scenarios on $2$ common-sense reasoning tasks and $4$ domain-specific question answering tasks without compromising accuracy.
@@ 2.2 Mixture of Experts (MoE)
LoRA Mixture-of-Experts (Li et al., 2024; Wu et al., 2024a) is a efficient approach to scale the number of parameters while maintaining the same computational bounds. LoRA MoE utilizes the top-k router to assign each token to the LoRA experts (Lepikhin et al., 2020). The router is a linear layer that maps the input hidden state $\mathbf{h}$ to a probability distribution of candidate experts.
Let $\mathbf{h}^{\ell}_{i}\in\mathbb{R}^{1\times d}$ $(1\leq i\leq s,1\leq\ell\leq L)$ denote the output hidden state of the $i$-th token at the $\ell$-th layer of the LLM, where $L$ is the number of LLM layers and $d$ is the hidden dimension. With $\mathbf{W}_{r}^{\ell}$ as the trainable router weight at layer $\ell$, the top-k gate router chooses $k$ experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
$R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-k}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (2)
@@
$\text{MixLoRA}(\mathbf{h}^{\ell})=\sum_{k=1}^{K}R^{\ell}(\mathbf{h}^{\ell})_{k}E^{\ell}_{k}(\mathbf{h}^{\ell}),\quad E^{\ell}_{k}(\mathbf{h}^{\ell})=\mathbf{W}^{\ell}\cdot\mathbf{h}^{\ell}+\mathbf{B}^{\ell}_{i}\mathbf{A}^{\ell}_{i}\cdot\mathbf{h}^{\ell}$ (3)
where $\mathbf{W}$ is the pretrained weights of the feed-forward network (FFN) layer and $\mathbf{B}^{\ell}_{i}\mathbf{A}^{\ell}_{i}$ is the i-th LoRA expert.
@@ 2.3 Aleatoric and Epistemic Uncertainties
Epistemic uncertainty is highly related to several well-known limitations of generative models. For example, it has been observed that when an LLM is pretrained on a diverse range of text data, it is generally well-calibrated, i.e. the predicted probability of the next token generally aligns with what is observed in real text. However, after fine-tuning or alignment with human preferences, the calibration error deteriorates (Zhao et al., 2021; Achiam et al., 2023a). A related phenomenon is forgetting, where the performance of a fine-tuned LLM diminishes on tasks outside the scope of the target downstream task (Lin et al., 2023; Luo et al., 2023).
@@ 3 Methodology
Note that we hereby quantify uncertainty as a function of the input prompt $x$, since the distribution of model $M$ conditions on $x$. We hence name the task “functional-level uncertainty quatification”.
@@ 3.1 Functional-level Epistemic Uncertainty
Figure 1: Mixture of Experts (MoE) architecture to capture and calibrate functional-level epistemic uncertainty. Experts $B_{1...N}A_{1...N}$ capture functional relationships in the data throughout fine-tuning, the weights $w_{1...N}|h$ quantify the uncertainty in selecting these functional bases conditioned on the input hidden state $h$, which is the semantic representation of the input token $x$. In the UQ4CT workflow, we align this uncertainty with predictive correctness. When the router makes a correct prediction, the loss reinforces this decision, thereby increasing the router confidence in its selection, which aligns with a lower epistemic uncertainty. Conversely, when the router makes an incorrect prediction, the loss penalizes this selection, potentially causing the router to distribute its probabilities more broadly across experts, which is indicative of higher epistemic uncertainty.
Motivated by the decomposition in Eq. (3) for least squares loss, we may consider a general distance $d$, and define epistemic uncertainty that characterizes the variation caused by model training procedure. Specifically, we focus on the variation introduced in the model fine-tuning stage of LLMs. Mathematically, given prompt $x$, we consider the following definition of epistemic uncertainty:
$\text{Epistemic Uncertainty}=\mathbb{E}_{M\sim P(M|x)}\mathbb{E}_{M^{\prime}\sim P(M^{\prime}|x)}\left[\|\mathbb{E}_{P(a|M,x)}[e(a)]-\mathbb{E}_{P(a^{\prime}|M^{\prime},x)}[e(a^{\prime})]\|^{2})\right].$ (5)
@@ 3.2 Quantifying FEU with LoRA MoE Framework
We quantify the functional-level epistemic uncertainty (FEU) with the MoE architecture, which is represented by the embedding $e(a)$ in Eq. (5). As shown in Figure 1, the LoRA experts $B_{1...N}A_{1...N}$ capture important functional relationships in the data during fine-tuning. We treat these functions represented by the LoRA experts as basis functions. Conditional on the input prompts, the more complex functional relationships that recursively map inputs to outputs are represented as linear combinations (mixture of experts) of the basis functions. Uncertainty quantification in this functional space reduces to quantifying the uncertainty of the weights over the basis functions. In particular, the weights $w_{1...N}|h$ from the top-k router dynamically selects these basis functions conditioned on the input hidden state $h$. The top-k weights that produce the final output hidden state quantify the functional level epistemic uncertainty in the function selection.
In this mixture of LoRA experts architecture, we follow the routing mechanisms of the MoE layers as in Eq. (2) and (3). Specifically, we employ top-2 gate routers, which chooses the $2$ experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
$R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-2}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (6)
@@
$\text{FEU}(x)=\frac{1}{s}\sum^{s}_{i=1}\left[\frac{1}{L}\sum^{L}_{\ell=1}R^{\ell}(\mathbf{h}^{\ell}_{i})\right]$ (7)

### v2

@@ Abstract
  1. Accurate uncertainty quantification of large language models (LLMs) provides credibility measure over their outputs. However, fine-tuned LLMs often struggle with overconfidence in uncertain predictions due to the limitations in the models’ ability to generalize with limited data. Existing parameter efficient fine-tuning (PEFT) uncertainty quantification methods for LLMs focus on post fine-tuning stage and fall short of calibrating epistemic uncertainty. To address these limitations, we propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT), which captures and calibrates epistemic uncertainty over the space of functions that map input prompts to outputs. We implement UQ4CT during the fine-tuning stage via a mixture-of-experts framework that hierarchically decomposes the functional space. We demonstrate that UQ4CT reduces Expected Calibration Error (ECE) by more than $25\%$ while maintaining high accuracy across $5$ benchmarks. Even under distribution shift, UQ4CT maintains superior ECE performance with high accuracy, showcasing improved generalizability.
@@ 1 Introduction
Quantifying credibility of the outputs has been one of the most important problems around large language models (LLMs) (Chang et al., 2024). In particular, fine-tuned LLMs often struggle with overconfidence in gauging the uncertain of their outputs—oftentimes due to limited training data—and does not reflect the true credibility of the answers that they generate (Xiao et al., 2022a; Zhou et al., 2022; He et al., 2023; Tian et al., 2023; OpenAI, 2023). Such overconfidence has become a crucial problem in safety-related decision making and in fields where data is relatively limited, such as formal proof generation, climate science, and healthcare (Singhal et al., 2022; Wu et al., 2023a; Lampinen et al., 2023; Li et al., 2022). Methods that enhance uncertainty quantification of fine-tuned LLMs are sought after for the assurance of trustworthy predictions in applications.
One salient challenge of uncertainty quantification in large models is the trade-off among accuracy, calibration, and tractability. One would first like to maintain the capability of the LLMs and not sacrifice accuracy of the generated outputs. Given this desideratum, a number of existing works focus on perturbing the prompts that are fed into the language models and quantify the uncertainty associated with such perturbations (Hou et al., 2023; Farquhar et al., 2024; Gao et al., 2024). These methods capture the aleatoric uncertainty assuming that the trained model corresponds to the ground-truth data distribution. However, these approaches cannot capture the epistemic uncertainty that arises from the model being trained on limited data. To capture this epistemic uncertainty, one would need to introduce changes or perturbations to the model in accordance with the data distribution. On the tractability side, it is not realistic to consider perturbations to all the model parameters for an LLM (Minaee et al., 2024).
We therefore propose Functional-Level Uncertainty Quantification for Calibrated Fine-Tuning (UQ4CT) to calibrate the epistemic uncertainty in LLMs. Since we only focus on the epistemic uncertainty over the functional space that maps input prompts to the outputs, we no longer need to consider uncertainty in the entire parameter space. Instead, we only need to measure the uncertainty regarding the functional relationships. Due to the immense size of the dataset and the transformer network, we assume that various compositions of parts of the pretrained network cover all the interesting functional relationships arising from the underlying data.
@@
In particular, we make use of the low rank adaptation (LoRA, Hu et al. (2021a)) ensemble at each layer of the LLM to form the basis functions. We then utilize the mixture of experts (MoE) architecture (Li et al., 2024) to hierarchically combine the basis functions and form the functional space. During fine-tuning, UQ4CT simultaneously learns the parameters in the LoRA experts and the prompt-dependent mixture weights to output the calibrated distribution over the functional space. During inference, the LoRA experts capture different functional relationships in the fine-tuning data throughout training, and the MoE routers dynamically select these functional bases conditioned on the input. The selection process models the functional level epistemic uncertainty, and consequently captures the uncertainty in the output space. We calibrate functional level epistemic uncertainty to align with predictive correctness during training time. This approach significantly improves uncertainty estimations of the model on its predictions without compromising accuracy. To summarize, our contributions include:
@@
  • A novel epistemic uncertainty quantification approach with MoE architecture during fine-tuning to quantify functional level epistemic uncertainty and align with predictive correctness, which mitigates overconfidence issue and improves generalizability.
@@
  • A novel calibration loss function incorporating predictive correctness to calibrate the prompt-dependent LoRA mixture for better uncertainty estimation.
@@
  • More than $25\%$ expected calibration error (ECE) reduction on $4$ common-sense reasoning tasks and $1$ domain-specific question answering task; improved ECE performance under distribution shifts on $2$ common-sense reasoning tasks and $4$ domain-specific question answering tasks without compromising accuracy.
@@ 2.2 Mixture of Experts (MoE)
LoRA Mixture-of-Experts (Li et al., 2024; Wu et al., 2024a) is an efficient approach to scale the number of parameters while maintaining the same computational bounds. LoRA MoE utilizes the top-k router to assign each token to the LoRA experts (Lepikhin et al., 2020). The router is a linear layer that maps the input hidden state $\mathbf{h}$ to a probability distribution of candidate experts.
The plain transformer block in a large language model consists of the $q,k,v$ encoding layers ($\text{FFN}_{q,k,v}$), layer norm (LN) and feed forward layer (FFN), together with residual connections. Formally, given $\mathbf{h}_{o}=[s_{0},s_{1},...,s_{n}]$ as the tokenized input text, the output of $\ell$-th layer is generated as:
$\mathbf{z}^{\ell}=f_{attn}(\text{FFN}_{q,k,v}(\text{LN}(\mathbf{h}^{\ell-1})))+\mathbf{h}^{\ell-1},\quad\mathbf{h}^{\ell}=\text{FFN}(\text{LN}(\mathbf{z}^{\ell}))+\mathbf{z}^{\ell}.$ (2)
Here, $f_{attn}$ represents the attention calculation step.
Let $\mathbf{h}^{\ell}_{i}\in\mathbb{R}^{1\times d}$ $(1\leq i\leq s,1\leq\ell\leq L)$ denote the output hidden state of the $i$-th token at the $\ell$-th layer of the LLM, where $L$ is the number of LLM layers and $d$ is the hidden dimension. With $\mathbf{W}_{r}^{\ell}$ as the trainable router weight at layer $\ell$, the top-n gate router $R(\cdot)$ chooses n experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
$R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-n}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (3)
@@
$\text{MixLoRA}(\mathbf{h}^{\ell})=\sum_{k=1}^{K}R^{\ell}(\mathbf{h}^{\ell})_{k}E^{\ell}_{k}(\mathbf{h}^{\ell}),\quad E^{\ell}_{k}(\mathbf{h}^{\ell})=\mathbf{W}^{\ell}\cdot\mathbf{h}^{\ell}+\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}\cdot\mathbf{h}^{\ell}$ (4)
where $\mathbf{W}$ is the pretrained weights of the target fine-tuning layer and $\mathbf{B}^{\ell}_{k}\mathbf{A}^{\ell}_{k}$ is the k-th LoRA expert.
With MixLoRA defined in Equation 4, we can apply MixLoRA layers at $q,k,v$ encoding and FFN layers:
$\mathbf{z}^{\ell}=f_{attn}(\text{MixLoRA}_{q,k,v}(\text{LN}(\mathbf{h}^{\ell-1})))+\mathbf{h}^{\ell-1},\quad\mathbf{h}^{\ell}=\text{MixLoRA}(\text{LN}(\mathbf{z}^{\ell}))+\mathbf{z}^{\ell}.$ (5)
Note that the calculation of $E^{\ell}_{k}$ involves $\mathbf{W}^{\ell}$, which represents the weight of the original ATTN or FFN layer.
@@ 2.3 Aleatoric and Epistemic Uncertainties
Epistemic uncertainty is highly related to several well-known limitations of generative models. For example, it has been observed that when an LLM is pretrained on a diverse range of text data, it is generally well-calibrated, i.e. the predicted probability of the next token generally aligns with what is observed in real text. However, after fine-tuning or alignment with human preferences, the calibration error deteriorates (Zhao et al., 2021; Achiam et al., 2023a). A related phenomenon is forgetting, where the performance of a fine-tuned LLM diminishes on tasks outside the scope of the target task (Lin et al., 2023; Luo et al., 2023).
@@ 3 Methodology
Note that we hereby quantify uncertainty as a function of the input prompt $x$, since the distribution of model $M$ conditions on $x$. We hence name the task “functional-level uncertainty quantification”.
@@ 3.1 Functional-level Epistemic Uncertainty
Figure 1: Mixture of Experts (MoE) architecture to capture and calibrate functional-level epistemic uncertainty. Experts $B_{1...K}A_{1...K}$ capture functional relationships in the data throughout fine-tuning, the weights $\alpha_{1...K}$ quantify the uncertainty in selecting these functional bases conditioned on the input hidden state $h$, which is the semantic representation of the input token $x$. In the UQ4CT workflow, we align this uncertainty with predictive correctness. When the router makes a correct prediction, the loss reinforces this decision, thereby increasing the router confidence in its selection, which aligns with a lower epistemic uncertainty. Conversely, when the router makes an incorrect prediction, the loss penalizes this selection, potentially causing the router to distribute its probabilities more broadly across experts, which is indicative of higher epistemic uncertainty.
Motivated by the decomposition in Eq. (3) for least squares loss, we consider a general distance $d$, and formally define epistemic uncertainty as the variance caused by model training procedure. Specifically, we focus on the variation introduced in the model fine-tuning stage of LLMs. Mathematically, given prompt $x$, we consider the following definition of epistemic uncertainty:
$\text{Epistemic Uncertainty}=\mathbb{E}_{M\sim P(\cdot|x)}\mathbb{E}_{M^{\prime}\sim P(\cdot|x)}\left[\|\mathbb{E}_{P(a|M,x)}[e(a)]-\mathbb{E}_{P(a^{\prime}|M^{\prime},x)}[e(a^{\prime})]\|^{2})\right]$ (7)
Here, $a^{\prime}$ and $a$ represents the outputs sampled from the parameter mixtures $M^{\prime}$ and $M$ conditioned on the prompt $x$. The epistemic uncertainty measures the least squares distance between $e(a)$ and $e(a^{\prime})$.
3.2 Decomposition of the Functional Space
Due to the immense size of the dataset and the transformer network, we assume that various compositions of parts of the pretrained network cover all the functional relationships in the underlying data. Then the functional space can be decomposed into a mixture of the LoRA ensembles. Naively, this decomposition writes:
$f(\mathbf{x})=\mathbf{h}^{L}=\sum_{k^{1}=1}^{K}\sum_{k^{2}=1}^{K}\cdots\sum_{k^{L}=1}^{K}\alpha_{k^{1},k^{2},\dots,k^{L}}g_{k^{L}}^{L}\left(g_{k^{L-1}}^{L-1}\left(\dots g_{k^{1}}^{1}(\mathbf{x})\right)\right),$ (8)
$g_{k^{\ell}}^{\ell}(\mathbf{h}^{\ell})=E_{k^{\ell}}^{\ell}\left(f_{trans}^{\ell}(\mathbf{h}^{\ell})\right),$ (9)
where $f_{trans}^{\ell}$ represents the necessary non-fine-tuned operations within a transformer block (i.e. layer norm, attention calculation, etc.). This linear decomposition requires examining $K^{L}$ parameters to span the entire functional space.
We instead consider a hierarchical decomposition of $f(\mathbf{x})$:
$f(\mathbf{x})=\mathbf{h}^{L}=\sum_{k^{L}=1}^{K}\alpha^{L}_{k^{L}}g_{k^{L}}^{L}\Bigg{(}\sum_{k^{L-1}=1}^{K}\alpha^{L-1}_{k^{L-1}}g_{k^{L-1}}^{L-1}\Bigg{(}\dots\sum_{k^{1}=1}^{K}\alpha^{1}_{k^{1}}g_{k^{1}}^{1}(\mathbf{x})\Bigg{)}\Bigg{)}.$
Here the output function is determined by the combinations of the weights $\alpha^{\ell}_{k^{\ell}}$ across all the layers. As a result, we only need $K\cdot L$ parameters to represent various compositions of the functional relationships.
To further improve the efficiency of representation, we apply a sparse dynamic routing and choose the mixture weights $\alpha_{k^{\ell}}=R^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell}))_{k^{\ell}}$ that depend on the input from the previous layer:
$\mathbf{h}^{\ell+1}=\sum_{k^{\ell}=1}^{K}R^{\ell}(f_{trans}^{\ell}(\mathbf{h}^{\ell}))_{k^{\ell}}g^{\ell}_{k^{\ell}}(\mathbf{h}^{\ell}).$ (10)
@@ 3.3 Quantifying FEU with LoRA MoE Framework
We quantify the functional-level epistemic uncertainty (FEU) with the MoE architecture, which is represented by the embedding $e(a)$ in Eq. (7). As shown in Figure 1, the LoRA experts that are made of matrices $B^{1...L}_{1...K},A^{1...L}_{1...K}$ in Eq. (1) capture important functional relationships in the data during fine-tuning. We treat these functional relationships represented by the LoRA experts as basis functions. As shown in Eq. (10), conditional on the input prompts, the more complex functional relationships that recursively map inputs to outputs are represented as linear combinations (mixture of experts) of the basis functions, hierarchically through each layer.
Uncertainty quantification in this functional space reduces to quantifying the uncertainty of the weights over the basis functions. As shown in Figure 1, the weights $\alpha_{k}^{\ell}$ from the top-n router dynamically mix these basis functions given the input hidden state. The top-n weights that produce the final output hidden state quantify the functional level epistemic uncertainty in the function selection. With mixture of LoRA experts architecture, we follow routing mechanisms of the MoE layers as in Eq. (3) and (4). Specifically, we employ top-2 gate routers, which chooses $2$ experts with highest probability given a hidden state $\mathbf{h}_{i}^{\ell}$:
$\vspace{-0.5mm}R^{\ell}(\mathbf{h}^{\ell}_{i})=\text{KeepTop-2}(\text{Softmax}(\mathbf{W}^{\ell}_{r}\cdot\mathbf{h}_{i}^{\ell})).$ (11)
@@
$\text{FEU}(x)=\frac{1}{s}\sum^{s}_{i=1}\left[\frac{1}{L}\sum^{L}_{\ell=1}R^{\ell}(\mathbf{h}^{\ell}_{i})\right].$ (12)